# 3D Gaussian Splatting as Markov Chain Monte Carlo

### v2

@@ 1 Introduction
However, despite the various extensions, a common shortcoming of these methods is that they mostly rely on the same initialization and densification strategy for placing the Gaussians: either the one originally suggested by [14], or other recently proposed variations [4]. Specifically, they rely on carefully engineered cloning and splitting heuristics for placing Gaussians [14, see “adaptive density control”]. Depending on the state of each Gaussian, they are cloned, split, or pruned, which is the primary way to control the number of Gaussians within the 3DGS representation. Moreover, Gaussians are regularly ‘reset’ by setting their opacities to small values to remove floaters in the representations. This heuristic-based approach requires multiple hyperparameters to be carefully tuned and, as we will show later on in the paper, can fail in some scenes.
@@
Given this view, the heuristics involved in the densification and pruning of Gaussians, as well as resetting their opacities, are no longer necessary. 3D Gaussians are simply the samples used for MCMC, thus exploring their sample locations is naturally dealt through SGLD updates. Any modification to the set of Gaussians, including increasing or decreasing their cardinality, can be reformulated as a deterministic state transition—relocation of Gaussians—where we move our sample (the set of Gaussians) to another sample (another set of Gaussians with different configuration). Importantly, to minimally disturb the MCMC sampling chain, we make sure that the two states, before and after the move, are of similar probability.  [^1] This implies that the training loss value does not change as we alter combinatorial properties, such as the number of Gaussians within the representation, preventing the training process from becoming unstable. This simple strategy, under the MCMC framework, is enough to provide renderings of high quality, beyond what is provided by the conventional heuristics.
In more detail, we propose to relocate Gaussians using a ‘cloning’ strategy, where we move ‘dead’ Gaussians (with low opacity) to where other ‘live’ Gaussians exist, but in a way that has minimal impact on the rendering, and thus on the probability distribution of the Gaussians. In other words, we set the composition of the cloned Gaussians to render the same images as before cloning.  [^2] While a modification of cloning [4] was recently suggested in a pre-print as a way to ensure the rendering is equal at the Gaussian centers, we show that this is not enough—the whole Gaussian must be considered. Without our careful strategy MCMC sampling provides sub-optimal training. Finally, to encourage efficient use of the Gaussians, we apply L1-regularization. As the extent of a Gaussian is defined both the opacity and the scale of Gaussians, we apply our regularization to both of them. This effectively encourages them to ‘disappear’ if unnecessary.
@@ 3.1 Brief review of 3D Gaussian Splatting
where $\mathcal{P}$ is the data-dependent probability density function for the distribution one wishes to sample from, and $\boldsymbol{\epsilon}$ is the noise distribution for exploration. The hyperparameters $a$ and $b$ together control the trade-off between convergence speed and exploration333 In typical SGLD formalism, $b$ is typically expressed as a function of $a$ and an additional hyper-parameter, but we rewrite it in this form without any loss of generality following [15]. We note the striking similarity between 4 and 5. In other words, by having the loss as the negative log likelihood of the underlying distribution,
@@
the equations become identical if $\lambda_{\text{lr}}{=}-a$ and $b{=}0$. Hence, the standard Gaussian Splatting optimization could be understood as having Gaussians that are sampled from a likelihood distribution that is tied to the rendering quality.
@@ 3.3 Updating with Stochastic Gradient Langevin Dynamics
In Equation 7, it is important that the noise term $\boldsymbol{\epsilon}$ is designed carefully. The noise term $\boldsymbol{\epsilon}$ needs to be added in a way that it can be ‘balanced’ by the gradient term $\nabla_{\mathbf{g}_{i}}\mathcal{L}_{\text{total}}$, or otherwise 7 reduces to random updates. For example, it is quite common for Gaussians to be narrow when reconstructing scenes to represent lines and edges. Should the added noise force Gaussians to move to locations outside of the previous support regions, this random walk would be irrecoverable, breaking the ergodicity required for MCMC. We further notice that exploration is not critical for opacity, scale, and color, and we do not add noise to these parameters—we empirically found that adding noise to them to slightly harm performance; see Section 4.1. Finally, as we are interested in the ‘converged’ quality not just exploration, we reduce the amount of noise when Gaussians are well-behaved, that is, when their opacities are high enough to be guided well by the gradients.
We thus design the noise term only on the center locations of the Gaussians such that it is dependent on their covariances and also its opacities, as well as the learning rate:
$\boldsymbol{\epsilon}=\lambda_{\text{lr}}\cdot{\sigma}\bigl{(}-k({o}-t)\bigr{)}\cdot\mathbf{\Sigma}\eta,$ (8)
where $\eta\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, ${\sigma}$ is the sigmoid function, and $k$ and $t$ are hyperparameters controlling the sharpness of the sigmoid, which we set as $k{=}100$ and $t{=}(1-0.005)$ to make a sharp transition function that goes from zero to one, centered around the default pruning threshold of 3D Gaussian Splatting [14] for the opacity values of Gaussians. In simple terms, 8 perturbs a Gaussian with anisotropic noise with the same anisotropy profile $\mathbf{\Sigma}$ of the Gaussian, while the sigmoid term reduces the effect of noise on opaque Gaussians.
@@ 3.4 Heuristics as state transitions via relocation
We now discuss how heuristics in 3D Gaussian Splatting can be rewritten as simple state transitions. In 3D Gaussian Splatting, heuristics are used to ‘move’, ‘split’, ‘clone’, ‘prune’, and ‘add’ Gaussians to encourage more ‘live’ Gaussians (${o}_{i}{\geq}0.005$). [^4] This is the default threshold used in 3D Gaussian Splatting [14]. We explain all of these modifications moving from one sample state $\mathbf{g}^{old}$ to another sample state $\mathbf{g}^{new}$. This applies also to cases where the number of Gaussians changes, as one could think of the state with a smaller number of Gaussians simply as the equivalent state with more Gaussians, but with those that have zero opacity, that is, dead Gaussians. Importantly, for these kinds of deterministic moves to be integrated into MCMC frameworks, it is important that they do not cause MCMC sampling to collapse. Specifically, we aim to preserve the probability of the sample state before and after the move that is, $\mathcal{P}(\mathbf{g}^{new})=\mathcal{P}(\mathbf{g}^{old})$ such that the move can be seen as simply hopping to another sample with equal probability. We now detail how we achieve this.
@@
While our method results in $\mathcal{P}(\mathbf{g}^{new})\approx\mathcal{P}(\mathbf{g}^{old})$, it is not exact. Hence we apply this move every 100 iterations to avoid disruptions in the training process. To choose where to move, for each dead Gaussian, we first chose a target Gaussian to move/teleport to via multinomial sampling of the live Gaussians with the probabilities proportional to their opacity values. Please note that only after all movement decisions have been made we apply 9. Finally, as we rely on the Adam optimizer, moment statistics should also be adjusted. We reset the moment statistics for the target Gaussian (the original one that is cloned) so that it is biased to stay stationary, while for the new ones (source) we retain the moment statistics to encourage exploration. [^5] This is because ‘dead’ (source) Gaussians are dominated by the noise term in 7, and the moment statistics are hence appropriate to foster exploration.
@@ 3.5 Encouraging fewer Gaussians
To make effective use of the memory and compute while improving the performance, we encourage Gaussians to disappear in non-useful locations and ‘respawn’. As the existence of a Gaussian is effectively determined by its opacity ${o}$ and covariance $\mathbf{\Sigma}$, we apply regularization to both of these. Our full training loss is:
@@ 4 Experiments
We use various datasets, both synthetic and real. Specifically, as in 3DGS [14], we use all scenes from NeRF Synthetic [22] dataset, the same two scenes used in [14] of Tank & Temples [17], and Deep Blending [13] and all publicly available scenes from MipNeRF 360 [2].  [^6] We do not use ‘Flower’ and ‘Treehill’ scenes from MipNeRF 360 [2] as they are not publicly available. We further use all scenes from the OMMO [21] dataset as in [9], for the large scenes with distant objects it provides. For MipNeRF 360 [2], to make our results compatible with [14], we downsample the indoor scenes by a factor of two, and the outdoor scenes four. For OMMO [21] scene #$\#01$, we downsample the images four times to keep the image size reasonable ($1000\times 750$). For all other scenes, we use the original image resolutions. In the main paper, we report our results by summarizing the average statistics for each dataset. Results for individual scenes, including the standard deviation of multiple runs, can be found in Appendix B. License information for each dataset can be found in Appendix F.
@@
We compare against conventional 3DGS [14], with both random and SfM point cloud-based initialization strategies—we denote the former with the ‘Random’ suffix. We use their official code for our experiments. We also report the original numbers in 3DGS [14], but where we correct their LPIPS scores, as reported by [4]. We also include state-of-the-art baselines for each dataset.
4.1 Results
Performance with the same number of Gaussians.
As the number of Gaussians is directly related to the quality of novel-view rendering, we first compare our method with existing baselines using the same number of Gaussians as 3DGS [14]. In more details, we simply set the number of Gaussians used in the original 3DGS [14] and set it as our maximum number of Gaussians to be used during training and inference. We show the quantitative results in Section 4.1, and qualitative highlights in Section 4. As shown, our method provides better performance to 3DGS [14]. Interestingly, our method, whether initialized randomly or with SfM point clouds, only displays minor variations in performance, thanks to the exploration that our MCMC formulation provides. This allows our method to outperform 3DGS [14], and regardless of the initialization strategy.
@@ 4.1 Results
3DGS [14$\rightarrow$ [4] 33.32 /    -    /    - 28.69 / 0.87 / 0.22 23.14 / 0.84 / 0.21 - -
3DGS [14] – Random 33.42 / 0.97 / 0.04 27.89 / 0.84 / 0.26 21.93 / 0.79 / 0.27 29.55 / 0.90 / 0.33 28.24 / 0.88 / 0.24
Ours – Random 33.80 / 0.97 / 0.04 29.72 / 0.89 / 0.19 24.21 / 0.86 / 0.19 29.71 / 0.90 / 0.32 29.31 / 0.90 / 0.20
3DGS [14] - 29.30 / 0.88 / 0.21 23.67 / 0.84 / 0.22 29.64 / 0.90 / 0.32 28.83 / 0.89 / 0.22
Ours - 29.89 / 0.90 / 0.19 24.29 / 0.86 / 0.19 29.67 / 0.89 / 0.32 29.52 / 0.91 / 0.20
@@
3DGS [14]– Random 27.89 / 0.84 / 0.26 22.72 / 0.75 / 0.34
Ours – Random 29.72 / 0.89 / 0.19 29.64 / 0.89 / 0.19
Sensitivity to initialization.
An important benefit of our method is that it allows exploration through the MCMC sampling scheme, removing the heavy reliance of 3DGS [14] on initialization. To verify this, we deviate from the default initialization strategy of 3DGS [14], which is to randomly place Gaussians within 3×3$\times$3 × the camera extent as defined by [14]. Instead, we use 1×1$\times$1 × the camera extent. We summarize our results for the average of all scenes in MipNeRF 360 [2] in Table 2. As shown, our method provides robustness to the initialization strategy, whereas 3DGS [14] requires careful initialization. This is also evident in Section 4.1, where our results with random initialization remain competitive with those initialized by SfM point clouds.
@@
Finally, we also explored various ways of adding noise to parameters other than the locations as well, but none worked better. We report one such attempt in Tab. 3, where we, in addition to our location noise in 8 apply $\mathcal{N}(0,1)$ to scale and rotation and $\mathcal{N}(0,0.1)$ for opacity, and decrease them exponentially with a decay rate of 0.9995 for each iteration. This slightly worse performance suggest that the other parameters do not require as much exploration as the locations do.

### v3

@@ 1 Introduction
Neural Radiance Fields (NeRF) have been at the forefront of these advancements, providing impressive results through implicit modeling via neural networks, the landscape is further evolving with the introduction of 3D Gaussian splatting.
@@
However, despite the various extensions, a common shortcoming of these methods is that they mostly rely on the same initialization and densification strategy for placing the Gaussians: either the one originally suggested by [19], or other recently proposed variations [4]. Specifically, they rely on carefully engineered cloning and splitting heuristics for placing Gaussians [19, see “adaptive density control”]. Depending on the state of each Gaussian, they are cloned, split, or pruned, which is the primary way to control the number of Gaussians within the 3DGS representation. Moreover, Gaussians are regularly ‘reset’ by setting their opacities to small values to remove floaters in the representations. This heuristic-based approach requires multiple hyperparameters to be carefully tuned and, as we will show later on in the paper, can fail in some scenes.
@@
Given this view, the heuristics involved in the densification and pruning of Gaussians, as well as resetting their opacities, are no longer necessary. 3D Gaussians are simply the samples used for MCMC, thus exploring their sample locations is naturally dealt through SGLD updates. Any modification to the set of Gaussians, including increasing or decreasing their cardinality, can be reformulated as a deterministic state transition—relocation of Gaussians—where we move our sample (the set of Gaussians) to another sample (another set of Gaussians with different configuration). Importantly, to minimally disturb the MCMC sampling chain, we make sure that the two states, before and after the move, are of similar probability. This implies that the training loss value does not change as we alter combinatorial properties, such as the number of Gaussians within the representation, preventing the training process from becoming unstable. This simple strategy, under the MCMC framework, is enough to provide renderings of high quality, beyond what is provided by the conventional heuristics.
In more detail, we propose to relocate Gaussians using a ‘cloning’ strategy, where we move ‘dead’ Gaussians (with low opacity) to where other ‘live’ Gaussians exist, but in a way that has minimal impact on the rendering, and thus on the probability distribution of the Gaussians. In other words, we set the composition of the cloned Gaussians to render the same images as before cloning. While a modification of cloning [4] was recently suggested as a way to ensure the rendering is equal at the Gaussian centers, we show that this is not enough—the whole Gaussian must be considered. Without our careful strategy MCMC sampling provides sub-optimal training. Finally, to encourage efficient use of the Gaussians, we apply L1-regularization. As the extent of a Gaussian is defined both the opacity and the scale of Gaussians, we apply our regularization to both of them. This effectively encourages them to ‘disappear’ if unnecessary.
@@ 3.1 Brief review of 3D Gaussian Splatting
where $\mathcal{P}$ is the data-dependent probability density function for the distribution one wishes to sample from, and $\boldsymbol{\epsilon}$ is the noise distribution for exploration. The hyperparameters $a$ and $b$ together control the trade-off between convergence speed and exploration111 In typical SGLD formalism, $b$ is typically expressed as a function of $a$ and an additional hyper-parameter, but we rewrite it in this form without any loss of generality following [20]. We note the striking similarity between 4 and 5. In other words, by having the loss as the negative log likelihood of the underlying distribution,
@@
the equations become identical if $\lambda_{\text{lr}}{=}-a$ and $b{=}0$. Hence, the standard Gaussian Splatting optimization could be understood as having Gaussians that are sampled from a likelihood distribution that is tied to the rendering quality. We further note that this addition of noise to optimization is highly related to traditional optimization methods that inject noise [35, 30, 8] or that perform perturbed gradient descent [17, 18]. Here, we formulate it as MCMC in favour of the probabilistic relationship that it provides in 6, and the removal of heuristics that it enables, which we discuss in Sec. 3.4.
@@ 3.3 Updating with Stochastic Gradient Langevin Dynamics
In Equation 7, it is important that the noise term $\boldsymbol{\epsilon}$ is designed carefully. The noise term $\boldsymbol{\epsilon}$ needs to be added in a way that it can be ‘balanced’ by the gradient term $\nabla_{\mathbf{g}_{i}}\mathcal{L}_{\text{total}}$, or otherwise 7 reduces to random updates. For example, it is quite common for Gaussians to be narrow when reconstructing scenes to represent lines and edges. Should the added noise force Gaussians to move to locations outside of the previous support regions, this random walk would be irrecoverable, breaking the MCMC sampling chain.
We further notice that exploration is not critical for opacity, scale, and color, and we do not add noise to these parameters. In fact, we empirically found that adding noise to them to slightly harm performance; see Section 4.1. Conversely, it has been shown by [11] that 3DGS reconstruction can be harmed significantly when areas of space are left unexplored, e.g. due to missing points in the initialization. This suggests potential for noise-based exploration to improve results for both random initialization as well as SFM-derived initializations which suffer from missing geometry.
Finally, as we are interested in the ‘converged’ quality not just exploration, we reduce the amount of noise when Gaussians are well-behaved, that is, when their opacities are high enough to be guided well by the gradients. Thus, we design the noise term only on the locations of the Gaussians such that it is dependent on their covariances and also its opacities, as well as the learning rate:
$\boldsymbol{\epsilon}_{\mu}=\lambda_{\text{lr}}\cdot{\sigma}\bigl{(}-k(t-{o})\bigr{)}\cdot\mathbf{\Sigma}\eta\quad\text{ where }\quad\boldsymbol{\epsilon}=[\boldsymbol{\epsilon}_{\mu},\mathbf{0}].$ (8)
where $\eta\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, ${\sigma}$ is the sigmoid function, and $k$ and $t$ are hyperparameters controlling the sharpness of the sigmoid, which we set as $k{=}100$ and $t{=}(0.005)$ to make a sharp transition function that goes from zero to one, centered around the default pruning threshold of 3D Gaussian Splatting [19] for the opacity values of Gaussians. In simple terms, 8 perturbs a Gaussian with anisotropic noise with the same anisotropy profile $\mathbf{\Sigma}$ of the Gaussian, while the sigmoid term reduces the effect of noise on opaque Gaussians.
@@ 3.4 Heuristics as state transitions via relocation
Inspired by jump and resampling moves in MCMC [26, 15], we now discuss how heuristics in 3D Gaussian Splatting can be rewritten as simple state transitions. In 3D Gaussian Splatting, heuristics are used to ‘move’, ‘split’, ‘clone’, ‘prune’, and ‘add’ Gaussians to encourage more ‘live’ Gaussians (${o}_{i}{\geq}0.005$). [^2] This is the default threshold used in 3D Gaussian Splatting [19]. We explain all of these modifications moving from one sample state $\mathbf{g}^{old}$ to another sample state $\mathbf{g}^{new}$. This applies also to cases where the number of Gaussians changes, as one could think of the state with a smaller number of Gaussians simply as the equivalent state with more Gaussians, but with those that have zero opacity, that is, dead Gaussians. Importantly, for these kinds of deterministic moves to be integrated into MCMC frameworks, it is important that they do not cause MCMC sampling to collapse. Specifically, we aim to preserve the probability of the sample state before and after the move that is, $\mathcal{P}(\mathbf{g}^{new})=\mathcal{P}(\mathbf{g}^{old})$ such that the move can be seen as simply hopping to another sample with equal probability. We now detail how we achieve this.
@@
While our method results in $\mathcal{P}(\mathbf{g}^{new})\approx\mathcal{P}(\mathbf{g}^{old})$, it is not exact. Hence we apply this move every 100 iterations to avoid disruptions in the training process. To choose where to move, for each dead Gaussian, we first chose a target Gaussian to move/teleport to via multinomial sampling of the live Gaussians with the probabilities proportional to their opacity values. Please note that only after all movement decisions have been made we apply 9. Finally, as we rely on the Adam optimizer, moment statistics should also be adjusted. We reset the moment statistics for the target Gaussian (the original one that is cloned) so that it is biased to stay stationary, while for the new ones (source) we retain the moment statistics to encourage exploration. This is because ‘dead’ (source) Gaussians are dominated by the noise term in 7, and the moment statistics are hence appropriate to foster exploration.
@@ 3.5 Encouraging fewer Gaussians
To make effective use of the memory and compute while improving the performance, we encourage Gaussians to disappear in non-useful locations and ‘respawn’ elsewhere. As the existence of a Gaussian is effectively determined by its opacity ${o}$ and covariance $\mathbf{\Sigma}$, we apply regularization to both of these. Our full training loss is:
@@ 4 Experiments
We use various datasets, both synthetic and real. Specifically, as in 3DGS [19], we use all scenes from NeRF Synthetic [28] dataset, the same two scenes used in [19] of Tank & Temples [22], and Deep Blending [16] and all publicly available scenes from MipNeRF 360 [2]. We do not use ‘Flower’ and ‘Treehill’ scenes from MipNeRF 360 [2] as they are not publicly available. We further use all scenes from the OMMO [27] dataset as in [10], for the large scenes with distant objects it provides. For MipNeRF 360 [2], to make our results compatible with [19], we downsample the indoor scenes by a factor of two, and the outdoor scenes four. For OMMO [27] scene #$\#01$, we downsample the images four times to keep the image size reasonable ($1000\times 750$). For all other scenes, we use the original image resolutions. In the main paper, we report our results by summarizing the average statistics for each dataset. Results for individual scenes, including the standard deviation of multiple runs, can be found in Appendix B. License information for each dataset can be found in Appendix E.
@@
We compare against conventional 3DGS [19], with both random and SfM point cloud-based initialization strategies. We use their official code for our experiments. We also report the original numbers in 3DGS [19], but where we correct their LPIPS scores, as reported by [4]. We also include state-of-the-art baselines for each dataset.
@@
3DGS [19$\rightarrow$ [4] 33.32 /    -    /    - 28.69 / 0.87 / 0.22 23.14 / 0.84 / 0.21 - -
3DGS [19] (Random) 33.42 / 0.97 / 0.04 27.89 / 0.84 / 0.26 21.93 / 0.79 / 0.27 29.55 / 0.90 / 0.33 28.24 / 0.88 / 0.24
Ours (Random) 33.80 / 0.97 / 0.04 29.72 / 0.89 / 0.19 24.21 / 0.86 / 0.19 29.71 / 0.90 / 0.32 29.31 / 0.90 / 0.20
3DGS [19]          (SfM) - 29.30 / 0.88 / 0.21 23.67 / 0.84 / 0.22 29.64 / 0.90 / 0.32 28.83 / 0.89 / 0.22
Ours (SfM) - 29.89 / 0.90 / 0.19 24.29 / 0.86 / 0.19 29.67 / 0.89 / 0.32 29.52 / 0.91 / 0.20
4.1 Results
Performance with the same number of Gaussians.
As the number of Gaussians is directly related to the quality of novel-view rendering, we first compare our method with existing baselines using the same number of Gaussians as 3DGS [19]. In more details, we simply set the number of Gaussians used in the original 3DGS [19] and set it as our maximum number of Gaussians to be used during training and inference. We show the quantitative results in Section 4, and qualitative highlights in Figure 2. As shown, our method provides better performance than 3DGS [19]. Interestingly, our method, whether initialized randomly or with SfM point clouds, only displays minor variations in performance, thanks to the exploration that our MCMC formulation provides. This allows our method to outperform 3DGS [19], and regardless of the initialization strategy.
@@ 4.1 Results
Sensitivity to initialization.
An important benefit of our method is that it allows exploration through the MCMC sampling scheme, removing the heavy reliance of 3DGS [19] on initialization. To verify this, we deviate from the default initialization strategy of 3DGS [19], which is to randomly place Gaussians within 3×3$\times$3 × the camera extent as defined by [19]. Instead, we use 1×1$\times$1 × the camera extent.
@@
3DGS [19]   (Random) 27.89 / 0.84 / 0.26 22.72 / 0.75 / 0.34
Ours (Random) 29.72 / 0.89 / 0.19 29.64 / 0.89 / 0.19
We summarize our results for the average of all scenes in MipNeRF 360 [2] in Table 2. As shown, our method provides robustness to the initialization strategy, whereas 3DGS [19] requires careful initialization. This is also evident in Section 4.1, where our results with random initialization remain competitive with those initialized by SfM point clouds.
@@
We also explore in Tab. 3 adding noise to other parameters. In addition to our location noise in 8, we apply $\mathcal{N}(\mathbf{0},\mathbf{I})$ to scale and rotation and $\mathcal{N}(\mathbf{0},0.1\times\mathbf{I})$ for opacity, and decrease them exponentially with a decay rate of 0.9995 for each iteration. This slightly worse performance suggests that the other parameters do not require as much exploration as the Gaussian locations do.
On Tank & Temples [22], we additionally investigate the design choice of our noise scheduler by removing the covariance term and the opacity term. Without the covariance term, we achieve a PSNR of 23.16, whereas with the covariance term 24.21. Without the opacity term, we achieve 22.47, which is also performing worse than with the two terms. We note that removing the opacity term requires $\lambda_{\text{noise}}$ to be set to 0.05 or the model fails to train (PSNR <{<}< 7).
Finally, we also compare the impact that noise scheduling has. Using the same exponential scheduling as in 3DGS [19]—which is what we use—achieves best performance of 24.21 PSNR, while a linear scheduler gives 17.64. We further compare with the scheduler suggested in [30], which gives 22.46.
@@
Table 4: Timings – We report the timings of our method with various opacity regularizer ($\lambda_{o}$) settings and the maximum number of Gaussians to 300k. Our method can deliver faster training while still maintaining reconstruction quality.
3DGS [19] Ours ($\lambda_{o}{=}0.01$) Ours ($\lambda_{o}{=}0.001$) Ours ($\lambda_{o}{=}0.01$, 300k)
PSNR 31.7 32.5 32.4 31.8
Time 25 minutes 42 minutes 30 minutes 21 minutes
4.2 Computational time
As our method results in the same 3D Gaussian Splat representation as prior work, inference time is the same with 3DGS [19], which is highly efficient. To provide an exact comparison, we took our Gaussians trained at 1M Gaussian, and re-measured the single optimization iteration time for our method and the original 3DGS. In this case, ours takes 80 milliseconds while 3DGS takes 76 milliseconds. That is, the added time for sampling and noise addition is not substantial, even with our implementation that implements resampling naively with PyTorch [33]’s torch.multinomial—this could be further accelerated with a CUDA implementation.
Furthermore, the configuration of Gaussians (i.e. where they are, their sizes and opacity) matters greatly to the runtime because they affect the speed of rasterization (among them the opacity regularizer affects the speed the greatest). Hence, we evaluate the runtime for the ‘Room’ scene in MipNeRF 360 dataset, all with SfM initializations and maximum number of 1.5M Gaussians per the original 3DGS implementation. We report the timings in Tab. 4. As shown, our method, while it may take longer to achieve the highest PSNR, trains faster when a similar PSNR is desired. An independent implementation of our method [45] also confirms a 20% reduction in training time and a 65% reduction in required memory when using our method.